OpenStudy (anonymous):

Check all polynomials that are perfect square trinomials: 9x2 − 6x + 1 36b2 − 24b + 8 16x2 + 24x + 9 4a2 − 10a + 25

OpenStudy (campbell_st):

use the discriminant \[\Delta = b^2 - 4ac\] when \[\Delta = 0\] the quadratic is a perfect square in the first question a = 9, b = -6 and c = 1 then \[\Delta = (-6)^2 - 4 \times9\times1 = 0\] so a perfect square

OpenStudy (anonymous):

that'll work or see if it looks like \(a^2+2ab+b^2\) because that is what you get when you square: \((a+b)^2=a^2+2ab+b^2\)

OpenStudy (anonymous):

im soo good at guessing lol